Please explain how to identify the acute and obtuse angle bisectors using the angle bisectory formula. 

Grade:11

1 Answers

Ajay Verma
9 years ago
if two lines are:
L1 : a1x + b1y + c1= 0
L2 : a2x + b2y + c2= 0


then eqn of angle bisectors
(a1x + b1y + c1) / {(a1)2 + (b1)2}1/2= +-(a2x + b2y + c2) / {(a2)^2+ (b2)^2}^1/2


we write the eqns of L1 and L2 in that way....that constent terms are positive.
now.. we check..

1. if a1*a2 + b1*b2 > 0
then for taking + sign in bisector formula we get Obtuse angle bisector..
for taking - sign in bisector formula we get acute angle bisector..

2. if a1*a2 + b1*b2 < 0
then for taking + sign in bisector formula we get acute angle bisector..
for taking - sign in bisector formula we get obtuse angle bisector..



thanks and regards.